OpenStudy (sammyalabamy):

Find the product of the complex number and its conjugate. 1 + 2i Select one: a. 1 - 4i b. -3 c. 5 d. 1 + 4i

OpenStudy (alexandervonhumboldt2):

the congugat of a+bi is a-bi. so the congugate of 1+2i is 1-2i. (1+2i)*(1-2i)=1*1-2i+2i-(2i*2i)=1-(2*-1)=1+2=3

OpenStudy (alexandervonhumboldt2):

woops: (1+2i)*(1-2i)=1-2i+2i+4=1+4=5

OpenStudy (alexandervonhumboldt2):

sorry made a mistake

OpenStudy (sammyalabamy):

Thank you!

OpenStudy (alexandervonhumboldt2):

np :)

OpenStudy (alexandervonhumboldt2):

:D :P you are welcome